OpenStudy (anonymous):

Which graph represents the solutions to the inequality |2x - 4| less than or greater to 14? number line with an open circle on negative 5, shading to the right and an open circle on 9, shading to the left number line with a closed circle on negative 5, shading to the right and a closed circle on 9, shading to the left number line with an open circle on negative 5, shading to the left and an open circle on 9, shading to the right number line with a closed circle on negative 5, shading to the left and a closed circle on 9, shading to the right

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

less than or greater or less than or equal?

OpenStudy (anonymous):

i dont know but this is the thing [2x - 4] ≤ 14

OpenStudy (anonymous):

The second choice \[ 2|x-2| \le 14\\ |x-2|\le 7\\ -5 \le x \le 9 \]

OpenStudy (anonymous):

thank you for all your help :D

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

could you help me with a couple more sorry?

OpenStudy (anonymous):

actually its just one more

OpenStudy (anonymous):

Post it as new problem and tag me.